Two rods each of mass m and length l  are joined to form cross. Moment of inertia passing through centre and perpendicular to plane is- Ml2/12, Ml2/6, Ml2/3, Ml2/2
 

Grade:

3 Answers

shashi K Sharma
9 years ago
Dear Student
 
The moment of inertia for each rod is I = ml2/12 , about the middle point and perpendicular to the plane thus making the sum total for two rods as 2I , i.e. ml2/6
Gagarin
5 years ago
2x1/12 ml^2 sin^2 45 = ml2/12So moment of inertia of the cross about bisector is I = ml^2/12
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

The moment of inertia for each rod is I = ml2/12 , about the middle point and perpendicular to the plane thus making the sum total for two rods as 2I , i.e. ml2/6

Thanks and Regards